Young Jeffrey is bored, and decides to throw some things out of the window for fun. But since he is also very curious, he decides to do so in a controlled way. Jeffrey takes a rock, his tennis ball, and one of his father\'s golf balls and sets them on the sill of an open window. Then he uses a flat board to carefully push them all out of the window at the same time. The 2nd story window is 10.0 feet above the ground, and all three objects hit the ground just 0.650 feet away from the base of the house at the same time, 0.788 seconds after they\'re dropped. The rock hits with a thud and stays put, but the other two objects bounce. After another 0.591 seconds, the tennis ball reaches a maximum height of 5.63 feet, while the golf ball is 7.88 feet off the ground and still rising. Both have travelled an additional 0.488 feet horizontally. Calculate the magnitude of the average velocity of each object during these 1.38 seconds. Which object has the largest instantaneous velocity at the end of this time period?
Answer #1

rock s1 = sqrt(10^2+0.65^2) = 10.02 feet


v1 = s1/t1 = 10.02/0.788 =12.71 ft/s


for tennis ball

x2 = 0.65+0.488 = 1.138 feet


y2 = 10-5.63 = 4.37 feet


s2 = sqrt(x2^2+y2^2) = 4.52 ft


v2 = s2/t2 = 4.52/1.38 = 3.27 ft/s


for golf


x3 = 0.65+0.488 = 1.138 feet

y3 = 10-7.88 = 2.12 ft

s3 = sqrt(x3^2+y3^2) = 2.42 ft


v3 = s3/t3 = 2.42/1.38 = 1.75 ft/s